OpenStudy (anonymous):

Solve M = 2x + 3y for y

mathslover (mathslover):

M = 2x + 3y subtract 2x from both sides wht r u getting ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

1x?

mathslover (mathslover):

M -2x = 2x + 3y - 2x right ?

OpenStudy (anonymous):

i suppose

mathslover (mathslover):

so what is 2x - 2x ?

OpenStudy (anonymous):

its 0? its nothing?

mathslover (mathslover):

right it is 0 so M - 2x = 0 + 3y hence M - 2x = 3y divide both sides by 3 wht r u getting ?

OpenStudy (anonymous):

1.5?

mathslover (mathslover):

\[\frac{M-2x}{3}=\frac{3y}{3}\] \[\frac{M-2x}{3}=\frac{\cancel{3}y}{\cancel{3}}\]

OpenStudy (anonymous):

how do you find the answer to M−2x over 3

mathslover (mathslover):

divided both sides by 3

OpenStudy (anonymous):

so what does it equal?